# The physical basis of teleological evolutions

### v11

@@ 3.2 Decohered teleological evolutions: what would they look like?
Decohered and endowed with quantum speedup, it is natural to wonder what the supposedly teleological evolutions of life would look like. All natural evolutions are necessarily described in a probabilistic way. We can think of the stochastic description of a Darwinian evolution: an alternation of random changes in the species DNA pool each followed by natural selection of the DNA changes more suited to the environment. Of course, in a classical description, transition probabilities can depend on the past of the process, not on the future of it. Instead, a teleological process must have an attractor in its future. This means that those DNA mutations that in the future will survive natural selection because more fit to the environment should have the transition toward them more probable than in the classical case.
Let us mention what could be an emblematic example of this kind of natural evolution. It is known that, for many millions of years, the ancestors of the flying birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave, and air-proof, feathers are of course ideal for flight. So, why anticipating something designed for flight millions of years in advance?
@@ 3.3 How to model teleological natural evolutions
Let us note that the evolution of life on Earth, obviously since the water stage phase of the latter, should have been almost completely independent of the evolution of the rest of the universe. If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious, approximate, way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact that it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[22\right]$.
@@ 3.5 Impact on philosophy
For short, let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view666 We can say that what generates the quantum superposition of the time-symmetric causal loops is the mutual causality between the initial and final measurement outcomes. and the causal view. The causal view would underlie the materialist, mechanistic and reductionist views, while the mutual causality view would underlie the idealist, teleological and monist views.
@@
Also Plato’s notion that ideas – fundamental laws – are objective perfections could have a correspondence with the Fine-tuned Universe version of the Anthropic Principle, the fact it asserts that the slightest change of the value of the fundamental constants would have produced a radically different universe most probably unable to host life.
(ii) The present notion of teleological evolution, namely of an evolution toward a goal with an attractor in the goal it will produce in the future, could also be seen in terms of a “systematically lucky” evolution, lucky from the standpoint of the objective of reaching that goal because attracted by it. Interestingly, also the notion of a systematically lucky evolution might have a precedent in pre-scientific thinking. It could be the Stoic notion of $\pi\rho o\nu o\iota\alpha$ (foreknowledge)$\left[31\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge. Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does in quantum algorithms. This notion of $\pi\rho o\nu o\iota\alpha$ would also fit well with the notion of teleological evolution understood as a stochastic processes with transition probabilities time-symmetrically affected by the past and future of the process, with the luckier transitions more probable than in the classical case.
@@ 4 Research prospects
The teleological view, with a present time-symmetrically affected by the past and the future, changes the way of viewing physical reality. Correspondingly, future work on the subject could have a broad spectrum.
@@
(ii) Look for toy models that capture the mechanism for which the mostly autonomous evolution of a subsystem (life on Earth), although decohered, inherits the quantum speedup and the consequent teleological character of the evolution of the quantum system (the universe). Note that, in this case, the interaction of the subsystem with the system plays a role opposite to that of decoherence; it is what enables the purely quantum evolution of the subsystem as an inseparable part of the quantum system. Also note that the same could apply to those aspects of the living subsystems suspected to host quantum features, like the chlorophyll cycle, orientation in migratory birds, and a quantum consciousness.
(iii) Look for a possible impact of the teleological view on quantum cosmology. According to current assumptions, even the “inanimate” part of the evolution of the universe should be an evolution toward life. According to the present quantum interpretation of the Anthropic principle, it could be endowed with quantum speedup and would consequently be teleological in character. The evolution of the universe could have an attractor in the future of it.
(iv) Look for toy models of teleological Darwinian evolutions. These evolutions could be either approximated by pure quantum processes endowed with quantum speedup or described as decohered stochastic processes with transition probabilities affected by both the past and the future of the process. Note that this might be essentially different from quantum algorithms.
(v) Look for applications of the teleological view to quantum self-organization. For example, let us consider abiogenesis. The jump from a soup of life precursor molecules to a self replicating system is still open research. Seeing this awkward classical jump as the inseparable part of an overall unitary evolution (that of the universe) with an attractor in the future full fledged self-replicating system might be appealing. One could look for toy models of this situation.
(vi) Finally, opposing schools of thought throughout the history of philosophy could be reviewed in the light of the opposition between the causal and mutual causality views, as exemplified in the previous subsection. The present theory of teleological evolutions might reduce the distance between the physical and philosophical views; there might be further room for cross-fertilization.
The above prospects are imagineable extensions of the present work. However, the notion that retrocausality becomes a legitimate direction of research in the quantum framework could have a deeper impact.
@@ 5 Positioning the work
The present retrocausal approach to the quantum speedup is naturally inspired by time-symmetric quantum mechanics and, of course, has in common with the attempts in question the fact of looking beyond the ordinary notion of causality, not in common the approach. In fact, these attempts explicitly rule out the possibility that quantum algorithms, with their causally ordered quantum circuits, can accommodate causality violations $\left[39,40\right]$. For us instead, with their quantum speedup over what is classically possible, they are the paradigmatic example of causality violation. The two approaches seem to address different kinds of causality violations. Possible complementarities should be the subject of further study.
@@
One is the celebrated EPR paper $\left[8\right]$. In it, Einstein, Podolsky, and Rosen already argued that the quantum description of quantum nonlocality is incomplete. As we have seen, the causal loops generated by time-symmetrizing the ordinary quantum description of quantum nonlocality play the role of the hidden variables they envisaged to physically (locally) explain it $\left[7\right]$.
Another precedent is the local, retrocausal, explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[43\right]$. Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems. Costa de Beauregard assumed that the outcome of the first of the two measurements (performed in the same basis) locally propagates backward in time, by the inverse of the unitary transformation that spatially separates the two subsystems, until these are not yet spatially separate; of course, this is equivalent to anticipate the measurement in question along a unitary space-separation process that does not change the measurement outcome, which is mathematically legitimate. There, always locally, it would change the state of the other observable in a way that ensures the appropriate correlation between the two future measurement outcomes.
@@
Interestingly, at the time, Costa de Beauregard was an assistant of De Broglie, who forbade him to publish that work for years until the publication, by Wheeler and Feynman, of their emitter-absorber theory that also resorted to retrocausality – see $\left[44\right]$ for an account of this. This tells us a lot about the position of physicists toward retrocausality, even at the top level. By the way, despite the importance of the author’s name, Costa de Beauregard’s explanation of quantum nonlocality never took off.
@@ 6 Conclusions
We provide the key points of the work in a consequential way.
We have shown that the sheer existence of a quantum computational speedup logically implies the mutually exclusive $or$ of well defined causal loops. In each of them, it is as if the problem-solver knew in advance, before beginning her problem-solving action, one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future. She can then use this knowledge to produce the solution with fewer computation steps than classically possible. The number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way benefitting of the advanced knowledge in question.
Of course, these causal loops must be implicit in the quantum algorithm that delivers the quantum speedup. Since the ordinary quantum description of quantum algorithm does not and, being causal in character, cannot describe them, we must conclude that it is incomplete. We have seen that the only possible way of completing it is by time-symmetrizing it.
We adopt an ordinary quantum description of quantum algorithms consisting of an initial measurement of the problem-setting in a quantum superposition or mixture of all the possible problem-settings, the unitary computation of the corresponding solution, and the final measurement of the latter.
We time-symmetrize this ordinary quantum description by requiring that the determination of the problem setting and the corresponding solution evenly shares between the initial measurement of the problem-setting and the final measurement of the solution. For each possible way of evenly sharing, the part of the selection ascribed to the initial measurement unitarily propagates forward in time until becoming the state immediately before the final measurement; that performed by the final measurement propagates backward in time, by the inverse of the time-forward unitary transformation, until replacing the previous initial measurement outcome. We should then take the quantum superposition of all these time-symmetrization instances.
This time-symmetrization leaves the unitary part of the ordinary quantum description of the quantum algorithm mathematically unaltered but changes the description of the behavior of causality along it. The single causal process of the ordinary quantum description is replaced by a quantum superposition of causal loops; as seen by the problem-solver (who should be concealed from the problem-setting selected by the initial measurement), these are the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
Since the time-symmetrization we are dealing with is the only possible way of obtaining a quantum description of the quantum algorithm that describes the causal loops that it must host, it is mandatory.
@@
In particular, this generalization covers the case of quantum nonlocality. The causal loops generated by time-symmetrizing its ordinary quantum description play the role of the hidden variables envisaged by Einstein and the others to physically (locally) explain it. Thus, the incompleteness we are currently ascribing to the ordinary quantum description of quantum algorithm is the same that Einstein and the others ascribed to the quantum description of quantum nonlocality. In both cases it is the incompleteness of the quantum description of a quantum correlation.
@@
We have seen that a quantum cosmological interpretation of the Fine-tuned Universe version of the Anthropic Principle shows that the evolution of the universe under the fundamental physical laws and constants builds up a quantum correlation between the value of the fundamental constants since the beginning of time and the “value” of life today (in particular, either “existence of sentient life” or “absence of life”). It should correspondingly be endowed with quantum speedup and consequently teleological in character. As we have seen, in the universe we are in this character would then be inherited by: (i) the evolution of the universe toward life, (ii) even if decohered, the evolution of life on Earth, and (iii) the evolutions of species.
We would conclude as follows. Explanations of physical phenomena involving retrocausality have been proposed in the past, in particular by Wheeler and Feynman $\left[46\right]$, Costa de Beauregard $\left[43\right]$, Aharonov, Cohen, and Elitzur $\left[36\right]$, and Cramer $\left[47\right]$. Intriguing as they are, it can be said that these explanations never took off as far as retrocausality never entered mainstream quantum mechanics. The present work makes a similar attempt in the new context of the quantum computational speedup. This should make a difference. In fact, we have shown that, necessarily, a quantum speedup with respect to what is classically possible logically implies the mutually exclusive $or$ of well defined causal loops. We have seen how consequently, in the complete quantum description of quantum algorithms, their single causal process should be replaced by a quantum superposition of causal loops. This should demonstrate that, contrary to current opinion, causal loops – of the (time-symmetric) kind we are dealing with – are commonplace in nonrelativistic quantum mechanics.
We have also seen that, under the quantum cosmological assumption that the universe evolves in a unitary quantum way, the same causal loops that explain the teleological character of quantum algorithms could be responsible for the teleological character of natural evolutions. This should falsify the assertion that a teleological character of the evolutions of the living has no possible physical basis.
@@ Acknowledgments
Thanks are due to Eliahu Cohen, Artur Ekert, Avshalom Elitzur and David Finkelstein for useful comments throughout the whole of the present retrocausal approach to the quantum speedup, Daniel Sheehan for organizing the San Diego AAAS Pacific Division conferences on retrocausality, a far-looking forum for the discussion of frontier, also unorthodox, physics, and Mario Rasetti for organizing the first series of international workshops on quantum communication and computation, the Elsag Bailey-ISI Turin Villa Gualino workshops of the years 1991-96.

### v12

@@ 3.2 Decohered teleological evolutions: what would they look like?
Decohered and endowed with quantum speedup, it is natural to wonder what the supposedly teleological evolutions of life would look like. All natural evolutions are necessarily described in a probabilistic way. The stochastic description of a Darwinian evolution would then be an alternation of random changes in the species DNA pool each followed by natural selection of the DNA changes more suited to the environment. Of course, in a classical description, transition probabilities can depend on the past of the process, not on the future of it. Instead, a teleological process must have an attractor in its future. This means that those DNA mutations that in the future will survive natural selection because more fit to the environment should have the transition toward them more probable than in the classical case.
Let us mention what would be an emblematic example of this kind of natural evolution:
It is known that, for many millions of years, the ancestors of the flying birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave, and air-proof, feathers are of course ideal for flight. So, why anticipating something designed for flight millions of years in advance?
@@ 3.3 How to model teleological natural evolutions
Let us note that the evolution of life on Earth should have been almost completely independent of the evolution of the rest of the universe. If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious, approximate, way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact that it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[22\right]$.
@@ 3.5 Impact on philosophy
For short, let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view777 We can say that what generates the quantum superposition of the time-symmetric causal loops is the mutual causality between the initial and final measurement outcomes. and the causal view. The causal view would underlie the materialist, mechanistic and reductionist views, while the mutual causality view would underlie the idealist, teleological and monist views.
@@
Also Plato’s notion that ideas – fundamental laws – are objective perfections could have a correspondence with the Fine-tuned Universe version of the Anthropic Principle, with the fact it asserts that the slightest change of the value of the fundamental constants would have produced a radically different universe most probably unable to host life.
(ii) The present notion of teleological evolution, namely of an evolution toward a goal with an attractor in the goal it will produce in the future, could also be seen in terms of a “systematically lucky” evolution, systematically lucky in reaching that goal because attracted by it. Interestingly, also the notion of a systematically lucky evolution might have a precedent in pre-scientific thinking. It could be the Stoic notion of $\pi\rho o\nu o\iota\alpha$ (foreknowledge)$\left[31\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge. Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does with quantum algorithms. This notion of $\pi\rho o\nu o\iota\alpha$ would also fit well with the notion of teleological evolution understood as a stochastic processes with transition probabilities time-symmetrically affected by the past and future of the process, with the luckier transitions more probable than in the classical case.
@@ 4 Research prospects
The teleological view, with a present time-symmetrically affected by the past and the future, changes the very way of viewing physical reality. Correspondingly, research prospects on the subject could have a broad spectrum.
@@
(ii) The advanced knowledge rule provides the number of oracle queries needed to solve in an optimal quantum way any oracle problem. The unifying character of this rule might suggest the possibility that there is a universal optimal quantum algorithm for this class of problems. Looking for this possibility could be another research prospect.
(iii) Look for toy models that capture the mechanism for which the mostly autonomous evolution of a subsystem (the evolution of life on Earth), although decohered, inherits the quantum speedup and the consequent teleological character of the evolution of the system (the evolution of a quantum universe toward life).
(iv) Look for a possible impact of the teleological view on quantum cosmology. According to current assumptions, also the “inanimate” part of the evolution of the universe should be an evolution toward life and would therefore have an attractor in its present state. Incidentally, since the assumption that the universe evolves in a unitary quantum way is essential to make a teleological character of natural evolutions possible, a possible empirical evidence of the teleological character of the evolutions of the living would also be empirical evidence of the quantum cosmological assumption.
(v) Look for toy models of teleological Darwinian evolutions. Note that these evolutions could be essentially different from quantum algorithms. For example, the fact that their teleological character should also be reflected in the functional structure of the living would have no parallel with quantum algorithms.
(vi) Look for applications of the teleological view to quantum self-organization. For example, one could look for toy models of the leap from a soup of life precursor molecules to a self-replicating system, a possibility considered in abiogenesis. Seeing this awkward (classical) leap as the part of a unitary evolution with an attractor in a future full fledged self-replicating system might be appealing.
(vii) Finally, opposing schools of thought throughout the history of philosophy could be reviewed in the light of the opposition between the causal and mutual causality views, as exemplified in the previous subsection. The fact that the philosophical notion of teleological evolution has a physical basis might reduce the distance between the physical and philosophical views; there might be further room for cross-fertilization.
The above prospects are imagineable extensions of the present work. However, the notion that the quantum framework can host a special form of retrocausality might have a broader impact.
@@ 5 Positioning the work
The present retrocausal approach to the quantum speedup is naturally inspired by time-symmetric quantum mechanics and, of course, has in common with the attempts in question the fact of looking beyond the ordinary notion of causality, not in common the approach. In fact, these attempts explicitly rule out the possibility that quantum algorithms, with their causally ordered quantum circuits, can accommodate causality violations $\left[39,40\right]$. For us instead, with their quantum speedup over what is classically possible, they are the paradigmatic example of causality violation. The two approaches seem to address different kinds of causality violations.
@@
One is the celebrated EPR paper $\left[3\right]$. In it, Einstein, Podolsky, and Rosen already argued that the quantum description of quantum nonlocality is incomplete. As we have seen, the causal loops generated by time-symmetrizing the ordinary quantum description of quantum nonlocality play the role of the hidden variables they envisaged to physically (locally) explain it $\left[8\right]$.
Another precedent is the local, retrocausal, explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[43\right]$. Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems. Costa de Beauregard assumed that the outcome of the first of the two measurements locally propagates backward in time, by the inverse of the unitary transformation that spatially separates the two subsystems, until these are not yet spatially separate. There, always locally, it would change the state of the other observable in a way that ensures the appropriate correlation between the two future measurement outcomes.
@@
Interestingly, at the time, Costa de Beauregard was an assistant of De Broglie, who forbade him to publish that work for years until the publication, by Wheeler and Feynman, of their emitter-absorber theory that – for the first time in the history of physics – resorted to retrocausality – see $\left[44\right]$ for an account of this. This tells us a lot about the position of physicists, even at the top level, toward retrocausality. By the way, despite the importance of the author’s name, Costa de Beauregard’s explanation of quantum nonlocality never took off.
@@ 6 Conclusions
We provide the line of thought followed in the work.
We have shown that the sheer existence of a quantum computational speedup logically implies the mutually exclusive $or$ of well defined causal loops. In each of them, it is as if the problem-solver knew in advance, before beginning her problem-solving action, one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future. She can then use this knowledge to produce the solution with fewer computation steps than classically possible. The number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way benefiting of the advanced knowledge in question.
Of course, these causal loops must be implicit in the quantum algorithm that delivers the quantum speedup. Since the ordinary quantum description of quantum algorithm does not and, being causal in character, cannot describe any causal loop, we must conclude that it is incomplete. We have seen that the only possible way of completing it is by time-symmetrizing it.
First, we should adopt an ordinary quantum description of quantum algorithms consisting of an initial measurement of the problem-setting in a quantum superposition or mixture of all the possible problem-settings, the unitary computation of the corresponding solution, and the final measurement of the latter.
Then, we time-symmetrize this ordinary quantum description by requiring that the determination of the problem-setting, and thus also that of the corresponding solution, evenly shares between the initial measurement of the problem-setting and the final measurement of the solution. For each possible way of evenly sharing, the part of the selection ascribed to the initial measurement unitarily propagates forward in time until becoming the state immediately before the final measurement; that performed by the final measurement propagates backward in time, by the inverse of the time-forward unitary transformation, until replacing the previous initial measurement outcome. We should then take the quantum superposition of all these time-symmetrization instances.
This time-symmetrization leaves the unitary part of the ordinary quantum description of the quantum algorithm mathematically unaltered but changes the description of the behavior of causality along it. The single causal process of the ordinary quantum description is replaced by a quantum superposition of causal loops; as seen by the problem-solver (who should be concealed from the part of problem-setting selected by the initial measurement), these are the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
Since the time-symmetrization we are dealing with turns out to be the only possible way of obtaining a quantum description of the quantum algorithm that describes the causal loops that it must host, it is mandatory.
@@
In particular, this generalization covers the case of quantum nonlocality. The causal loops generated by time-symmetrizing its ordinary quantum description play the role of the hidden variables envisaged by Einstein and the others to physically (locally) explain it. Thus, the incompleteness we are currently ascribing to the ordinary quantum description of quantum algorithm is exactly the same that Einstein and the others ascribed to the quantum description of quantum nonlocality. In both cases it is the incompleteness of the quantum description of a quantum correlation.
@@
We have seen that a quantum cosmological interpretation of the Fine-tuned Universe version of the Anthropic Principle shows that the evolution of the universe under the fundamental physical laws and constants builds up a quantum correlation between the value of the fundamental constants since the beginning of time and the “value” of life today (in particular, “existence of sentient life” or “absence of life”). It can correspondingly be endowed with quantum speedup and consequently be teleological in character. As we have seen, in the universe we are in, this possibility would be inherited by: (i) the evolution of the universe toward life, (ii) even if decohered, the evolution of life on Earth, and (iii) the evolutions of species. Under the conservative assumption that the search for the value of life by the evolution of the universe under the fundamental physical law and constants is random, this possibility of quantum speedup becomes an actual quantum speedup.
We would conclude as follows:
Explanations of physical phenomena involving retrocausality have been proposed in the past, in particular by Wheeler and Feynman $\left[46\right]$, Costa de Beauregard $\left[43\right]$, Aharonov, Cohen, and Elitzur $\left[36\right]$, and Cramer $\left[47\right]$. Intriguing as they are, it can be said that these explanations never took off, as no form of retrocausality ever entered mainstream quantum mechanics.
The present work makes a similar attempt in the new context of the quantum computational speedup. This should make a difference: the key fact is that any quantum speedup with respect to what is classically possible logically implies the mutual exclusive $or$ of well defined causal loops whose quantum superposition must be implicit in the quantum algorithm that delivers it. This also highlights the teleological character of quantum algorithms, their being evolutions toward a goal (the solution) with an attractor in the very goal they will produce in the future (again the solution).
Under the quantum cosmological assumption that the universe evolves in a unitary quantum way, the same causal loops that highlight the teleological character of quantum algorithms could be responsible for the teleological character of natural evolutions. Quantum cosmology is of course hypothetical; in any case, the assertion that a teleological character of the evolutions of the living has no possible physical basis would be falsified.
@@ Acknowledgments
Thanks are due to Eliahu Cohen, Artur Ekert, Avshalom Elitzur and David Finkelstein for useful comments throughout the whole of the present retrocausal approach to the quantum speedup, to Daniel Sheehan for organizing the San Diego AAAS Pacific Division conferences on retrocausality, a far-looking forum for the discussion of frontier, also unorthodox, physics, and to Mario Rasetti for organizing the first series of international workshops on quantum communication and computation, the Elsag Bailey-ISI Turin Villa Gualino workshops of the years 1991-96 that brought to the discovery of the major quantum algorithms.